The population N(t) of a species of micro-organism in a laboratory setting at any time t is established to vary under the influence of a certain chemical at a rate given by dN/dt=t-2et show that N(t)=t-2et+c.hence if N(0)=400 determine the population when t=5





1
Expert's answer
2021-07-06T14:47:13-0400
"\dfrac{dN}{dt}=t-2e^t"

Then


"dN=(t-2e^t)dt"

Integrate both sides


"\int dN=\int (t-2e^t)dt"

"N(t)=\dfrac{1}{2}t^2-2e^t+c"

Given "N(0)=400"


"400=\dfrac{1}{2}(0)^2-2e^0+c=>c=402"

"N(t)=\dfrac{1}{2}t^2-2e^t+402"


"N(5)=\dfrac{1}{2}(5)^2-2e^5+402=118"